OpenStudy (anonymous):

maybe im having a bad day, but im working with instantaneous velocity. how does 6(2+h)^2+5(2+h)+3=6h^2+29h+37? idk if im doing order of operations wrong or not, but i cannot get that answer.

OpenStudy (anonymous):

use BEDMAS, Brackets Exponents, division and multiplication, addition, subtraction

OpenStudy (anonymous):

why are you taking a derivative by first principles?

OpenStudy (anonymous):

=6 ( 4+4h +h^2) + 10 + 5h + 3 = 24 + 10 + 3 + 24h + 5h + 6h^2 = 6h^2 +29h +37

OpenStudy (anonymous):

thank you. i knew i was messing up the exponent part of the problem